# On dense orbits in the space of subequivalence relations

### v2

@@ 3.3 Seeing that the space of subequivalence relations is closed
This implies that the space of graphings on $\mathcal{R}$ (symmetric subrelations of $\mathcal{R}$) is closed in $\mathrm{MAlg}(\mathcal{R},M)$, because it is the set of fixed points of $\sigma$. In particular, we recover the following fact from [Kec, Chap. 18].
@@
The space of graphings is Polish for the topology induced by the measure algebra $\mathrm{MAlg}(\mathcal{R},M)$.∎
@@
This could also be seen through the same approach as for Theorem 2.6, noting that the liminf of any sequence of graphings is a graphing.
@@
has thus measure at most $\epsilon/2$. We conclude that the set of $x\in\mathrm{dom}\;\varphi_{0}\psi_{0}$ such that $\varphi\psi(x)=\varphi_{0}\psi_{0}(x)$ contains $\mathrm{dom}\;\varphi_{0}\psi_{0}$ up to an $\epsilon/2+\epsilon/2=\epsilon$ error as wanted. The fact that the continuity also holds for the $m$-lower topology is then a direct consequence of Lemma 3.8 since elements of $[[\mathcal{R}]]$ have $M$-measure at most $1$. ∎
@@
Fixing a countable group $\Gamma$ such that $\mathcal{R}=\mathcal{R}_{\Gamma}$, the desired continuity follows from the previous lemma along with the formula
@@
We already observed that being reflexive exactly means containing $\Delta_{X}$ and thus is a closed condition by Lemma 3.13. The fact that symmetry is a closed condition is a reformulation of the fact that graphings form a closed set (Proposition 3.14). Finally transitivity is a closed condition by the previous proposition. So $\operatorname{\mathrm{Sub}}(\mathcal{R})$ is the intersection of three closed subsets, hence closed itself. ∎
@@
The map which takes a sequence $(\mathcal{R}_{n})$ to the equivalence relation $\bigvee_{n}\mathcal{R}_{n}$ it generates is continuous in the lower topology, in particular it is Baire class one for the strong topology.
@@ 4.1 Preliminaries on conditional measures and index
In order to work in the p.m.p. setup, we have to understand which sets can be taken to other sets by (pseudo) full group elements up to measure zero. This is done here through the concept of conditional measures, a low-tech version of the ergodic decomposition which was already present in Dye’s founding paper [Dye59]. We already used them in a previous paper on non-ergodic p.m.p. equivalence relations, see [LM16, Sec. 2], but we recently gave a more detailed exposition for general full groups in a joint work with Slutsky which we use here as a reference (see Appendix D in [LMS23]).
@@
Let $\mathcal{R}$ be a p.m.p. equivalence relation on $(X,\mu)$, denote by $M_{\mathcal{R}}$ the closed subalgebra of $(X,\mu)$ consisting of $[\mathcal{R}]$-invariant subsets, and by $\mathbb{E}_{\mathcal{R}}$ the projection $\mathrm{L}^{2}(X,\mu)\to\mathrm{L}^{2}(X,M_{\mathcal{R}},\mu)$. Given $A\in\mathrm{MAlg}(X,\mu)$, its $M_{\mathcal{R}}$-conditional measure $\mu_{\mathcal{R}}$ is the $M_{\mathcal{R}}$-measurable function
@@
It can be checked that $\mu_{\mathcal{R}}$ takes values in $[0,1]$, satisfies the usual axioms of a measure, and that elements of $[\mathcal{R}]$ preserve $\mu_{\mathcal{R}}$ (see [LMS23, Prop. D.6]). Say that $\mathcal{R}$ is ergodic when $M_{\mathcal{R}}=\{\emptyset,X\}$, then $\mathcal{R}$ is ergodic iff $\mu_{\mathcal{R}}=\mu$. By [LMS23, Prop. D.10], we moreover have:
@@
Proposition 4.3 (Maharam’s Lemma, see [LMS23, Thm. D.12]).
@@
  • has finite index in $\mathcal{R}$ if for almost all $x\in X$, the $\mathcal{R}$-equivalence class of $x$ is the reunion of a finite set of $\mathcal{S}$-classes.
@@ 4.2 Approximating the diagonal
    the closure of the $[\mathcal{R}]$-orbit of $\mathcal{S}$ contains $\Delta_{X}$;
@@
The implication (ii) $\implies$ (iii) is clear since $[\mathcal{R}]\leqslant\mathrm{Aut}(\mathcal{R})$.
@@
This inequality defines a closed set of subequivalence relations, but it is not satisfied by $\Delta_{X}$, which finishes the proof of (iii)$\Rightarrow$(i).
We finally prove that (i) implies ii. We fix $\mathcal{S}\in\operatorname{\mathrm{Sub}}(\mathcal{R})$ everywhere of infinite index. Let us also fix a uniquely generating sequence $(\varphi_{k})$ of moving partial involutions of $\mathcal{R}$ as provided by Proposition 3.4. Applying Proposition 3.1 to the sequence of $M$-finite measure subsets $\varphi_{k}$, it suffices to show that given some $k\in\mathbb{N}$ and $\epsilon>0$, we can find $T\in[\mathcal{R}]$ such that for all $i\in\{1,\dots,k\}$,
@@
Our construction makes a crucial use of a result of Eisenmann and Glasner: letting $\Gamma=\{\gamma_{n}\colon n\in\mathbb{N}\}$ be a dense subgroup of $[\mathcal{R}]$, by [EG16, Prop. 1.19] we have that $\Gamma$ acts highly transitively on the orbit of almost every $x\in X$, and we may as well restrict ourselves to the set of all such $x$’s.
We can now start building $T$ inductively as the increasing union of elements $\psi_{n}$ of the pseudo full group defined as follows.
First, $\psi_{0}$ is the restriction of $\gamma_{0}$ to the set of all $x\in X$ such that for all $i\in\{1,\dots,k\}$, we have $(\gamma_{0}x,\gamma_{0}\varphi_{i}(x))\not\in\mathcal{S}$. Then, assuming $\psi_{n}$ has been built, we extend it as $\psi_{n+1}$ by letting $\psi_{n+1}(x)=\gamma_{n+1}x$ if $x\not\in\mathrm{dom}\;\psi_{n}$ and $\gamma_{n+1}x\not\in\operatorname{\mathrm{rng}}\psi_{n}$ and for all $i\in\{1,\dots,k\}$
  • if $\psi_{n}\varphi_{i}(x)$ is defined then $(\gamma_{n+1}x,\psi_{n}\varphi_{i}(x))\not\in\mathcal{S}$.
  • if $\psi_{n}\varphi_{i}(x)$ is not defined then $(\gamma_{n+1}x,\gamma_{n+1}\varphi_{i}(x))\not\in\mathcal{S}$
Let $\psi=\bigcup_{n}\psi_{n}$, we need to argue that $\psi$ has full domain. Assume not, observe first that for almost all $x$ in the complement of the domain of $\psi$, the $\mathcal{R}$-class of $x$ intersects of the complement of the range of $\psi$, because otherwise we have a positive measure $\mathcal{R}$-invariant set such that $\psi^{-1}$ takes it into a subset of itself of smaller measure, contradicting that $\mathcal{R}$ is measure-preserving.
Now take $x\in X\setminus\mathrm{dom}\;\psi$ as above. Let $i_{1},...,i_{l}\in\{1,\dots,k\}$ be the indices such that $\psi\varphi_{i_{1}}(x),\dots,\psi\varphi_{i_{l}}(x)$ is defined, denote by $j_{1}\dots j_{k-l}$ the remaining indices. Since $\mathcal{S}$ has infinite index in $X\setminus\operatorname{\mathrm{rng}}\psi$ there is $z\in[x]_{\mathcal{R}}\cap X\setminus\operatorname{\mathrm{rng}}\psi$ and $z_{1},\dots,z_{k-l}\in[x]_{\mathcal{R}}\cap X\setminus\operatorname{\mathrm{rng}}\psi$ such that
  • for all $m\in\{1,\dots,l\}$ we have $(z,\psi_{n}\varphi_{i_{m}}(x))\not\in\mathcal{S}$ and
@@
By high transitivity, there are infinitely many $\gamma\in\Gamma$ such that
@@
This shows that almost all $x\in X\setminus\mathrm{dom}\;\varphi$ there is $n\in\mathbb{N}$ such that $\gamma_{n+1}x\not\in\operatorname{\mathrm{rng}}\psi$ and for all $i\in\{1,\dots,k\}$, we have :
  • if $\psi\varphi_{i}(x)$ is defined then $(\gamma_{n+1}x,\psi\varphi_{i}(x))\not\in\mathcal{S}$, and
  • if $\psi\varphi_{i}(x)$ is not defined then $(\gamma_{n+1}x,\gamma_{n+1}\varphi_{i}(x))\not\in\mathcal{S}$.
Let $n$ be the first integer such that the set of $x\in X\setminus\mathrm{dom}\;\varphi$ satisfying the above two conditions is non null. By the definition of $\psi$, this set should be contained in the domain of $\psi_{n+1}$ and hence of $\psi$, a contradiction.
So $\psi$ is everywhere defined; we just need to check that the full group element $T=\psi$ is as wanted. Let $x\in X$, take $n\geqslant-1$ such that $T(x)$ was defined at stage $n+1$ as $T(x)=\psi_{n+1}(x)=\gamma_{n+1}x$.
Let $i\in\{1,\dots,k\}$. If $\psi_{n}\varphi_{i}(x)$ was already defined, then $(\psi_{n}\varphi_{i}(x),\gamma_{n+1}x)\not\in\mathcal{S}$ by construction, which means that $(T\varphi_{i}(x),T(x))\not\in\mathcal{S}$. If $\psi_{n}\varphi_{i}(x)$ was not defined, there are two possibilites:
  • Either we defined at the same time $\psi_{n+1}(\varphi_{i}(x))=\gamma_{n+1}\varphi_{i}(x)$, in which case the second condition $(\gamma_{n+1}x,\gamma_{n+1}\varphi_{i}(x))\not\in\mathcal{S}$ guarantees $(\psi_{n}\varphi_{i}(x),\gamma_{n+1}x)\not\in\mathcal{S}$ and hence $(T\varphi_{i}(x),T(x))\not\in\mathcal{S}$.
  • Or we defined $\psi_{m+1}(\varphi_{i}(x))=\gamma_{m+1}(\varphi_{i}(x))$ at a later stager $m>n$, but then the first condition $(\gamma_{m+1}\varphi_{i}(x),\psi_{m}\varphi_{i}\varphi_{i}(x))\not\in\mathcal{S}$ and the fact that $\varphi_{i}$ is involutive imply that $(\psi_{m+1}\varphi_{i}(x),\psi_{m}(x))\not\in\mathcal{S}$ and hence $(T\varphi_{i}(x),T(x))\not\in\mathcal{S}$.
Since in all cases we reached the desired conclusion $(T\varphi_{i}(x),T(x))\not\in\mathcal{S}$, the proof is finished. ∎
@@ 4.3 Dense orbits in the space of hyperfinite subequivalence relations
By definition a CBER is called finite when all its equivalence classes are finite.
@@
Let $\mathcal{R}$ be a CBER, say that $\mathcal{R}$ is hyperfinite if it can be written as an increasing union of finite Borel subequivalence relations.
In the measured context, hyperfiniteness can be characterized in full group terms as follows: $\mathcal{R}$ is hyperfinite iff for all $T_{1},\dots,T_{n}\in[\mathcal{R}]$ and $\epsilon>0$, after throwing away a set $X^{\prime}$ of measure $\epsilon$, the equivalence relation generated by the restrictions $T_{1\restriction X\setminus X^{\prime}},...,T_{n\restriction X\setminus X^{\prime}}$ is finite.
@@
The above definition is Dye’s original notion of approximate finiteness for full groups [Dye59] (see also [LM14, Prop. 1.57] or [KM04, Lem. 10.4] for a proof of the aforementionned equivalence).
@@
[…]
@@
Let $\mathcal{R}$ be an ergodic p.m.p. equivalence relation. Then two finite subequivalence relations $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ of $\mathcal{R}$ are in the same $[\mathcal{R}]$-orbit iff for all $n\in\mathbb{N}$
@@
Our assumption then becomes: the equality $\mu(X^{1}_{n})=\mu(X^{2}_{n})$ holds for every $n\geqslant 1$. Let $<$ be a Borel linear order on $X$. We can then define for all $i\in\{1,2\}$ and $j\in\{1,\dots,n\}$ the Borel set
@@
Let $\mathcal{R}$ be a p.m.p. ergodic equivalence relation, let $\mathcal{S}\in\operatorname{\mathrm{Sub}}(\mathcal{R})$ be aperiodic and have everywhere infinite index. Then the closure of the $[\mathcal{R}]$-orbit of $\mathcal{S}$ contains $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$.
@@
It follows from the definition of hyperfiniteness that finite equivalence relations are dense in $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$, so it suffices to approximate every finite subequivalence relation of $\mathcal{R}$ by an element of the $[\mathcal{R}]$-orbit of $S$. Let $\mathcal{R}_{0}$ be such a finite equivalence relation. For each $n$ let
@@
Because $\mathcal{S}$ is aperiodic, we can find an element of the $[\mathcal{R}]$-orbit of $\mathcal{R}_{0}$ contained in $\mathcal{S}$. Indeed, by Maharam’s lemma we first have a partition of $X$ into pieces $(Y_{j,n})_{1\leqslant j\leqslant n}$ such that $\mu_{\mathcal{S}}(Y_{j,n})=\frac{\mu(X_{n})}{n}$. We then have for every $2\leqslant j\leqslant n$ some $\varphi_{j,n}\in[[\mathcal{S}]]$ such that $\mathrm{dom}\;\varphi_{j,n}=Y_{1,n}$ and $\operatorname{\mathrm{rng}}\varphi_{j,n}=Y_{j,n}$.
@@
By [IKT09, Cor. 5.4 (ii)], if $\mathcal{R}$ is aperiodic and comes from a measure-preserving action of a property (T) countable group, then there are no dense orbits in the space of subequivalence relations because any $\mathcal{S}$ with a dense orbit would have to have finite index in some restriction of $\mathcal{R}$, and hence cannot contain $\Delta_{X}$ in its orbit by Theorem 4.7. The general fact is that there cannot be dense orbits in the space of subequivalence relations of $\mathcal{R}$ as soon as $\mathcal{R}$ is not approximable as defined by Gaboriau and Tucker-Drob in [GT16]. Indeed if $\mathcal{S}$ has a dense orbit Lemma 2.2 provides a sequence $\mathcal{S}_{n}$ such that =lim $\mathcal{R}=\liminf\mathcal{S}_{n}$, while non approximability forces some restriction of $\bigcap_{n\geqslant N}\mathcal{S}_{n}$ to coincide with $\mathcal{R}$ on a positive measure subset $A$, contradicting the density of the orbit of $\mathcal{S}$ by Theorem C. For more examples of non approximable equivalence relations, the reader can consult the paper of Gaboriau and Tucker-Drob, where they obtain for instance a quantitative version of non-approximability for some equivalence relations coming from actions of product groups (see [GT16, Thm. 2.4]).
@@
    $\mathcal{S}$ is aperiodic and the closure of the $[\mathcal{R}]$-orbit of $\mathcal{S}$ contains $\Delta_{X}$;
@@
    The closure of the $[\mathcal{R}]$-orbit of $\mathcal{S}$ contains $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$;

### v3

@@ 3.3 Seeing that the space of subequivalence relations is closed
This implies that the space of (Borel) subgraphs of $\mathcal{R}$ (symmetric subsets of $\mathcal{R}\setminus\mathrm{id}_{X}$) is closed in $\mathrm{MAlg}(\mathcal{R},M)$, because it is the set of fixed points of $\sigma$ contained in $\mathcal{R}\setminus\mathrm{id}_{X}$. In particular, we recover the following fact from [Kec, Chap. 18].
@@
The space of subgraphs is Polish for the topology induced by the measure algebra $\mathrm{MAlg}(\mathcal{R},M)$.∎
@@
This could also be seen through the same approach as for Theorem 2.6, noting that the liminf of any sequence of subgraphs is a subgraph.
@@
has thus measure at most $\epsilon/2$. We conclude that the set of $x\in\mathrm{dom}\;\varphi_{0}\psi_{0}$ such that $\varphi\psi(x)=\varphi_{0}\psi_{0}(x)$ contains $\mathrm{dom}\;\varphi_{0}\psi_{0}$ up to an $\epsilon/2+\epsilon/2=\epsilon$ error as wanted. The fact that the continuity also holds for the $m$-lower topology is then a direct consequence of Lemma 3.8 since elements of $\mathcal{R}$]] have $M$-measure at most $1$. ∎
@@
Fixing a countable subgroup $\Gamma\leqslant[\mathcal{R}]$ such that $\mathcal{R}=\mathcal{R}_{\Gamma}$, the desired continuity follows from the previous lemma along with the formula
@@
We already observed that being reflexive exactly means containing $\Delta_{X}$ and thus is a closed condition by Lemma 3.13. The fact that symmetry is a closed condition is a reformulation of the fact that subgraphs form a closed set (Proposition 3.14). Finally transitivity is a closed condition by the previous proposition. So $\operatorname{\mathrm{Sub}}(\mathcal{R})$ is the intersection of three closed subsets, hence closed itself. ∎
@@
The map which takes a sequence $(\mathcal{R}_{n})$ of subequivalence relations of $\mathcal{R}$ to the equivalence relation $\bigvee_{n}\mathcal{R}_{n}$ they generate is continuous in the lower topology, in particular it is Baire class one for the strong topology.
@@ 4.1 Preliminaries on conditional measures and index
In order to work in the p.m.p. setup, we have to understand which sets can be taken to other sets by (pseudo) full group elements up to measure zero. This is done here through the concept of conditional measures, a low-tech version of the ergodic decomposition which was already present in Dye’s founding paper [Dye59]. We already used them in a previous paper on non-ergodic p.m.p. equivalence relations, see [LM16, Sec. 2], but we recently gave a more detailed exposition for general full groups in a joint work with Slutsky which we use here as a reference (see Appendix D in [LMS25]).
@@
Let $\mathcal{R}$ be a p.m.p. equivalence relation on $(X,\mu)$, denote by $M_{\mathcal{R}}$ the closed subalgebra of $(X,\mu)$ consisting of $\mathcal{R}$]-invariant subsets, and by $\mathbb{E}_{\mathcal{R}}$ the projection $\mathrm{L}^{2}(X,\mu)\to\mathrm{L}^{2}(X,M_{\mathcal{R}},\mu)$. Given $A\in\mathrm{MAlg}(X,\mu)$, its $M_{\mathcal{R}}$-conditional measure $\mu_{\mathcal{R}}$ is the $M_{\mathcal{R}}$-measurable function
@@
It can be checked that $\mu_{\mathcal{R}}$ takes values in $[0,1]$, satisfies the usual axioms of a measure, and that elements of $\mathcal{R}$] preserve $\mu_{\mathcal{R}}$ (see [LMS25, Prop. D.6]). Say that $\mathcal{R}$ is ergodic when $M_{\mathcal{R}}=\{\emptyset,X\}$, then $\mathcal{R}$ is ergodic iff $\mu_{\mathcal{R}}=\mu$. By [LMS25, Prop. D.10], we moreover have:
@@
Proposition 4.3 (Maharam’s Lemma, see [LMS25, Thm. D.12]).
@@
  • has finite index in $\mathcal{R}$ if for almost all $x\in X$, the $\mathcal{R}$-equivalence class of $x$ is the union of a finite set of $\mathcal{S}$-classes.
@@ 4.2 Approximating the diagonal
    the closure of the $\mathcal{R}$]-orbit of $\mathcal{S}$ contains $\Delta_{X}$;
@@
The implication (ii) $\implies$ (iii) is clear since $\mathcal{R}]\leqslant\mathrm{Aut}(\mathcal{R})$.
@@
[…]
We finally prove that (i) implies (ii). We fix $\mathcal{S}\in\operatorname{\mathrm{Sub}}(\mathcal{R})$ everywhere of infinite index. Let us also fix a uniquely generating sequence $(\varphi_{k})$ of moving partial involutions of $\mathcal{R}$ as provided by Proposition 3.4. Applying Proposition 3.1 to the sequence of $M$-finite measure subsets $\varphi_{k}$, it suffices to show that given some $k\in\mathbb{N}$ and $\epsilon>0$, we can find $T\in[\mathcal{R}]$ such that for all $i\in\{1,\dots,k\}$,
@@
Our construction makes a crucial use of a result of Eisenmann and Glasner: letting $\Gamma=\{\gamma_{n}\colon n\in\mathbb{N}\}$ be a dense subgroup of $\mathcal{R}$], by [EG16, Prop. 1.19] we have that $\Gamma$ acts highly transitively on the orbit of almost every $x\in X$, and we may as well restrict ourselves to the set of all such $x$’s. By the definition of high transitivity, this means that for all $x\in X$, every partial bijection between finite subsets of $\Gamma x$ is the restriction of some $\gamma\in\Gamma$.
We will now build $T$ as the increasing union of elements $\psi_{n}$ of the pseudo full group defined inductively as follows.
First, $\psi_{0}$ is the restriction of $\gamma_{0}$ to the set of all $x\in X$ such that for all $i\in\{1,\dots,k\}$, we have $(\gamma_{0}x,\gamma_{0}\varphi_{i}(x))\not\in\mathcal{S}$. Then, assuming $\psi_{n}$ has been built, we extend it as $\psi_{n+1}$ by letting
$\psi_{n+1}(x)=\gamma_{n+1}x$
if $x\notin\mathrm{dom}\;\psi_{n}$ and $\gamma_{n+1}x\notin\operatorname{\mathrm{rng}}\psi_{n}$ and for all $i\in\{1,\dots,k\}$
  1. (a)
    if $\varphi_{i}(x)\in\mathrm{dom}\;\psi_{n}$ then $(\gamma_{n+1}x,\psi_{n}\varphi_{i}(x))\not\in\mathcal{S}$;
  2. (b)
    if $\varphi_{i}(x)\notin\mathrm{dom}\;\psi_{n}$ then $(\gamma_{n+1}x,\gamma_{n+1}\varphi_{i}(x))\not\in\mathcal{S}$.
Let $\psi=\bigcup_{n}\psi_{n}$, we now prove the following central claim.
Claim.
The element of the pseudo full group $\psi$ has full domain.
Proof of the claim.
Assume not, observe first that for almost all $x$ in the complement of the domain of $\psi$, the $\mathcal{R}$-class of $x$ intersects of the complement of the range of $\psi$. Indeed otherwise we have a positive measure $\mathcal{R}$-invariant set such that $\psi^{-1}$ takes it into a subset of itself of smaller measure, contradicting that $\mathcal{R}$ is measure-preserving.
Now take $x\in X\setminus\mathrm{dom}\;\psi$ as above. Let $i_{1},...,i_{l}\in\{1,\dots,k\}$ be the indices such that $\psi\varphi_{i_{1}}(x),\dots,\psi\varphi_{i_{l}}(x)$ is defined, denote by $j_{1}\dots j_{k-l}$ the remaining indices. Since $\mathcal{S}$ has infinite index in $X\setminus\operatorname{\mathrm{rng}}\psi$, there are $z\in[x]_{\mathcal{R}}\cap X\setminus\operatorname{\mathrm{rng}}\psi$ and pairwise distinct $z_{1},\dots,z_{k-l}\in[x]_{\mathcal{R}}\cap X\setminus\operatorname{\mathrm{rng}}\psi$ such that
  • for all $m\in\{1,\dots,l\}$ we have $(z,\psi\varphi_{i_{m}}(x))\not\in\mathcal{S}$;
@@
By high transitivity, there is $\gamma\in\Gamma$ such that
@@
This shows that almost all $x\in X\setminus\mathrm{dom}\;\psi$ there is $n\in\mathbb{N}$ such that $\gamma_{n+1}x\not\in\operatorname{\mathrm{rng}}\psi$ and for all $i\in\{1,\dots,k\}$, we have :
  1. (a’)
    if $\varphi_{i}(x)\in\mathrm{dom}\;\psi$ then $(\gamma_{n+1}x,\psi\varphi_{i}(x))\not\in\mathcal{S}$;
  2. (b’)
    if $\varphi_{i}(x)\notin\mathrm{dom}\;\psi$ then $(\gamma_{n+1}x,\gamma_{n+1}\varphi_{i}(x))\not\in\mathcal{S}$.
Let $n$ be the first integer such that the set of $x\in X\setminus\mathrm{dom}\;\varphi$ satisfying the above conditions is non null. By the definition of $\psi$, this set should be contained in the domain of $\psi_{n+1}$ and hence of $\psi$, a contradiction.∎
So $\psi$ is everywhere defined; since $\mathcal{R}$ is measure-preserving this implies that it belongs to the full group, and we thus rather write it as $T=\psi$. We finally check that $T$ is as wanted.
Let $x\in X$, define $n\geqslant-1$ as the least integer such that $x\in\mathrm{dom}\;\psi_{n+1}$, so that $T(x)=\psi_{n+1}(x)=\gamma_{n+1}(x)$ by construction. Let $i\in\{1,\dots,k\}$.
If $\varphi_{i}(x)\in\mathrm{dom}\;\psi_{n}$, then $(\gamma_{n+1}x,\psi_{n}\varphi_{i}(x))\not\in\mathcal{S}$ by Condition (a), which means that $(T(x),T\varphi_{i}(x))\not\in\mathcal{S}$.
If $\varphi_{i}(x)\notin\mathrm{dom}\;\psi_{n}$, let $m$ be the least integer such that $\varphi_{i}(x)\in\mathrm{dom}\;\psi_{m+1}$. Then $m\geqslant n$ and $\psi_{m+1}\varphi_{i}(x)=\gamma_{m+1}\varphi_{i}(x)$ by construction. There are two possibilites:
  • Either $m=n$, then Condition (b) guarantees $(\gamma_{n+1}x,\gamma_{n+1}\varphi_{i}(x))\not\in\mathcal{S}$ so that $(\gamma_{n+1}x,\psi_{n+1}\varphi_{i}(x))\not\in\mathcal{S}$ and hence $(T(x),T\varphi_{i}(x))\not\in\mathcal{S}$.
  • Or $m>n$, but then $(\gamma_{m+1}\varphi_{i}(x),\psi_{m}\varphi_{i}\varphi_{i}(x))\not\in\mathcal{S}$ by Condition (a) applied to $x^{\prime}=\varphi_{i}(x)$ and $n^{\prime}=m$. The fact that $\varphi_{i}$ is involutive implies that $(\psi_{m+1}\varphi_{i}(x),\psi_{m}(x))\not\in\mathcal{S}$ and hence $(T\varphi_{i}(x),T(x))\notin\mathcal{S}$, which is equivalent to $(T(x),T\varphi_{i}(x))\not\in\mathcal{S}$.
Since in all cases we reached the desired conclusion $(T(x),T\varphi_{i}(x))\not\in\mathcal{S}$, the proof is finished. ∎
@@ 4.3 Dense orbits in the space of hyperfinite subequivalence relations
By definition a CBER is called finite when all its equivalence classes are finite, and hyperfinite if it can be written as an increasing union of finite Borel subequivalence relations. In our measured context, we ignore null sets and thus use the following definition.
@@
A non-singular equivalence relation $\mathcal{R}$ on $(X,\mu)$ is called hyperfinite when it admits a restriction to a full measure set which is hyperfinite in the above sense.
Hyperfiniteness can then be characterized in full group terms as follows: $\mathcal{R}$ is hyperfinite iff for all $T_{1},\dots,T_{n}\in[\mathcal{R}]$ and $\epsilon>0$, after throwing away a set $X^{\prime}$ of measure $\epsilon$, the equivalence relation generated by the restrictions $T_{1\restriction X\setminus X^{\prime}},...,T_{n\restriction X\setminus X^{\prime}}$ is finite.
@@
The above characterization of hyperfiniteness is Dye’s original notion of approximate finiteness for full groups [Dye59] (see also [LM14, Prop. 1.57] or [KM04, Lem. 10.4] for a proof of this characterization).
@@
Using $\liminf$’s as in the proof of Theorem 2.6 and the fact that the class of non-singular hyperfinite equivalence relations is stable under countable increasing unions, one can show that that $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$ is closed in $\operatorname{\mathrm{Sub}}(\mathcal{R})$ (see [Kec, Thm. 8.1]). Let us observe that this can also be seen as a consequence of the characterization via approximate finiteness: if $\mathcal{S}$ is not hyperfinite as witnessed by some $\epsilon>0$ and $T_{1},...,T_{n}$ its in full group, any $\mathcal{S}^{\prime}$ which contains the $T_{i}$’s in its full group up to an $\epsilon/2n$ error will also fail to be hyperfinite.
@@
Let $\mathcal{R}$ be an ergodic p.m.p. equivalence relation. Then two finite subequivalence relations $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ of $\mathcal{R}$ are in the same $\mathcal{R}$]-orbit iff for all $n\in\mathbb{N}$
@@
Our assumption then becomes: the equality $\mu(X^{1}_{n})=\mu(X^{2}_{n})$ holds for every $n\geqslant 1$. Let << be a Borel linear order on $X$. We can then define for all $i\in\{1,2\}$ and $j\in\{1,\dots,n\}$ the Borel set
@@
Let $\mathcal{R}$ be a p.m.p. ergodic equivalence relation, let $\mathcal{S}\in\operatorname{\mathrm{Sub}}(\mathcal{R})$ be aperiodic and have everywhere infinite index. Then the closure of the $\mathcal{R}$]-orbit of $\mathcal{S}$ contains $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$.
@@
It follows from the definition of hyperfiniteness that finite equivalence relations are dense in $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$, so it suffices to approximate every finite subequivalence relation of $\mathcal{R}$ by an element of the $\mathcal{R}$]-orbit of $S$. Let $\mathcal{R}_{0}$ be such a finite equivalence relation. For each $n$ let
@@
Because $\mathcal{S}$ is aperiodic, we can find an element of the $\mathcal{R}$]-orbit of $\mathcal{R}_{0}$ contained in $\mathcal{S}$. Indeed, by Maharam’s lemma we first have a partition of $X$ into pieces $(Y_{j,n})_{1\leqslant j\leqslant n}$ such that $\mu_{\mathcal{S}}(Y_{j,n})=\frac{\mu(X_{n})}{n}$. We then have for every $2\leqslant j\leqslant n$ some $\varphi_{j,n}\in[[\mathcal{S}]]$ such that $\mathrm{dom}\;\varphi_{j,n}=Y_{1,n}$ and $\operatorname{\mathrm{rng}}\varphi_{j,n}=Y_{j,n}$.
@@
By [IKT09, Cor. 5.4 (ii)], if $\mathcal{R}$ is aperiodic and comes from a measure-preserving action of a property (T) countable group, then there are no dense orbits in the space of subequivalence relations because any $\mathcal{S}$ with a dense orbit would have to have finite index in some restriction of $\mathcal{R}$, and hence cannot contain $\Delta_{X}$ in its orbit by Theorem 4.7. The general fact is that there cannot be dense orbits in the space of subequivalence relations of $\mathcal{R}$ as soon as $\mathcal{R}$ is not approximable as defined by Gaboriau and Tucker-Drob in [GT16]. Indeed if $\mathcal{S}$ has a dense orbit Lemma 2.2 provides a sequence $\mathcal{S}_{n}$ such that $\mathcal{R}=\liminf\mathcal{S}_{n}$, while non approximability forces some restriction of $\bigcap_{n\geqslant N}\mathcal{S}_{n}$ to coincide with $\mathcal{R}$ on a positive measure subset $A$, contradicting the density of the orbit of $\mathcal{S}$ by Theorem C. For more examples of non approximable equivalence relations, the reader can consult the paper of Gaboriau and Tucker-Drob, where they obtain for instance a quantitative version of non-approximability for some equivalence relations coming from actions of product groups (see [GT16, Thm. 2.4]).
@@
    $\mathcal{S}$ is aperiodic and the closure of the $\mathcal{R}$]-orbit of $\mathcal{S}$ contains $\Delta_{X}$;
@@
    The closure of the $\mathcal{R}$]-orbit of $\mathcal{S}$ contains $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$;